OpenStudy (jake124):

using rounding to estimate the product 4 3/4 x 7 2/17 A. about 28 B. about 32 C. about 35 D. about 40 MY ANSWER

OpenStudy (jake124):

@DallasCowboys88

OpenStudy (jake124):

@AloneS

alones (alones):

\[4\frac{ 3 }{ 4 } * 7\frac{ 2 }{ 17}\] same as \[\ 5*7=35\]

alones (alones):

I'd go with C.

OpenStudy (jake124):

thank you

alones (alones):

You welcome.